# JOGS: Joint Optimization of Pose Estimation and 3D Gaussian Splatting

### v1

@@ 4.2 Comparing with Baseline
Figure 2: Qualitative results of several representative samples picked from LLFF-NeRF, Tanks and Temples, Shiny. Our method achieves consistently high rendering quality across all scenes.
(a) GT (Ground Truth)
@@
Figure 3: The comparison of all the methods in scene details. Obviously, our method is better in the detail and texture of novel view synthesis due to the addition of camera pose optimization during training.
Table 4: Pose estimation performance comparison on LLFF dataset. Our method outperforms both COLMAP-free baseline methods (CFGS and GSHT). As discussed in Table˜2, results for fortress and leaves are omitted.
@@
Table 5: Ablation study of joint optimization across three benchmark datasets. Init means working with only pose initialization without iterative refinement, while Full means working with the full version of our method containing joint optimization of Gaussian points and camera poses. Best results are highlighted in bold.
[…]
Init Full Init Full Init Full
LLFF-NeRF 25.25 25.39 0.81 0.81 0.20 0.20
Tanks and Temples 25.94 26.91 0.86 0.88 0.14 0.13
Shiny 24.98 25.58 0.80 0.85 0.23 0.21
Mean 25.39 25.96 0.82 0.85 0.19 0.18
Novel View Synthesis Evaluation. As shown in Tables˜1, 2 and 3, both CFGS and GSHT suffer degraded reconstruction quality, primarily due to their reliance on temporal continuity—when pose changes become large (e.g., under frame-subsampling on Tanks and Temples), their results collapse, as shown by the sharp drop in PSNR. In contrast, our method is sequence-agnostic and remains robust even under aggressive subsampling. As illustrated in Figure˜2, our method generates sharper geometric features and more coherent textures, in contrast to the blurred reconstructions of CFGS and the fragmented surfaces of GSHT. Beyond holistic visual assessment, Figure˜3 presents fine-grained comparisons of structural details. Our method achieves superior fidelity in geometric preservation and texture reconstruction compared to baseline approaches.
In addition, it is noteworthy that in Figure˜2, the CFGS method produces noticeably blurred novel view synthesis images due to its inaccurate pose estimation. This issue becomes particularly pronounced in the detailed regions as illustrated in Figure˜3. As demonstrated in the figure, both CFGS and GSHT exhibit variations in scale and positional displacement of the display units. As shown in Figure˜3 (the first row), 3DGS shows obvious blurring around high-frequency structures such as the edge of the display. This is primarily due to the lack of joint camera pose optimization during training, where even slight pose inaccuracies can be amplified during dense rendering, leading to structural blur and color artifacts. We evaluate on the Shiny dataset, which features strong reflections and refractions. As shown in Table˜3, JOGS matches COLMAP+3DGS in overall metrics and significantly outperforms both COLMAP-free baselines.
Camera Pose Estimation. In Table˜4, we provide a quantitative comparison of camera pose estimation on the LLFF dataset. The estimated camera poses are first aligned in scale with the ground truth, following the alignment strategy proposed in GSHT, and evaluated in terms of ATE and RPE. As shown in Figure˜4, our method produces significantly more accurate results than both CFGS and GSHT, demonstrating its effectiveness in reducing both relative and absolute pose errors.
CFGS
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GSHT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fern Flower Orchids Room Trex Horns
Figure 4: Trajectory comparison of different methods across several scenes
4.3 Ablation Study
To validate the necessity of our joint optimization framework, we conduct an ablation study comparing two variants: (1) Initialization-only (using initialized poses without iterative refinement during training) and (2) Full method (with alternating Gaussian and pose optimization). Experiments are conducted on three benchmarks: Tanks and Temples, LLFF-NeRF and Shiny.
As shown in Table˜5, the full method outperforms the reduced initialization-only variant across all datasets. The ablation study demonstrates that our combined optimization framework effectively mitigates error accumulation and enhances the synthesis accuracy of novel view scenes by alternately updating Gaussian points and refining camera poses using LK3D.
5 Conclusion
In this paper, we introduce a novel view synthesis framework that jointly optimize pose estimation and 3D Gaussian splatting, without requiring camera poses as inputs. This framework outperforms state-of-the-art methods in both pose estimation accuracy and rendering quality, particularly under challenging conditions, by leveraging an alternating optimization strategy for 3D Gaussian representations and camera poses.
Appendix
Appendix A Additional experiments
Ballroom Refer to caption Refer to caption Refer to caption
Barn Refer to caption Refer to caption Refer to caption
Church Refer to caption Refer to caption Refer to caption
Family Refer to caption Refer to caption Refer to caption
CFGS GSHT Ours
Figure 5: Trajectory comparison on Ballroom, Barn, Church, and Family from the Tanks and Temples dataset.
Francis Refer to caption Refer to caption Refer to caption
Horse Refer to caption Refer to caption Refer to caption
Ignatius Refer to caption Refer to caption Refer to caption
Museum Refer to caption Refer to caption Refer to caption
CFGS GSHT Ours
Figure 6: Trajectory comparison on Francis, Horse, Ignatius, and Museum from the Tanks and Temples dataset.
@@ Appendix A Additional experiments
Table 6: Pose estimation performance comparison on Tanks and Temples dataset.
@@
Appendix B Optimization Strategy of Rotation Matrix.
Specifically, the rotation matrix $R$ is parameterized as:
$R=R_{z}(\gamma)R_{y}(\beta)R_{x}(\alpha),$ (11)
Appendix C Dataset licenses.
We use the following datasets:

### v2

@@ IV-A Datasets
Figure 2: Qualitative results of several representative samples from LLFF-NeRF, Tanks and Temples, Shiny. Our method achieves consistently high rendering quality across all scenes.
TABLE III: Quantitative comparison on Shiny. The original dataset contains eight scenes, but both CFGS and GSHT suffer from running errors and fail to report the final results in some scenes. Thus we only report the experimental results for four scenes.
Scene PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS CFGS GSHT Ours 3DGS CFGS GSHT Ours 3DGS CFGS GSHT Ours
Cd 28.29 26.60 26.44 28.18 0.94 0.87 0.90 0.94 0.12 0.17 0.16 0.12
Giants 21.69 14.37 16.01 20.52 0.72 0.45 0.36 0.68 0.28 0.47 0.62 0.25
Lab 28.54 26.26 27.99 29.28 0.93 0.82 0.91 0.94 0.15 0.18 0.15 0.15
Tools 24.59 12.44 11.67 24.33 0.84 0.50 0.47 0.83 0.35 0.59 0.60 0.32
Mean 25.77 18.51 19.05 25.58 0.86 0.58 0.59 0.85 0.23 0.39 0.43 0.21
(a) GT
Refer to captionRefer to caption
@@
Figure 3: Comparison of fine-grained scene details. Our method preserves sharper textures in novel-view synthesis, benefiting from camera pose optimization during training.
IV-B Evaluation Metrics
We employed the same evaluation metrics as CFGS [9] and GSHT [42]. For novel view synthesis, we used standard metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [48] and learned perceptual image patch similarity (LPIPS) [49]. For pose evaluation, we treated COLMAP-estimated poses as ground truth and measured absolute trajectory error (ATE), which includes relative rotation error (RPEr) and relative translation error (RPEt), along with relative pose error (RPE). ATE quantifies the discrepancy between estimated camera positions and ground truth, while RPE measures relative pose errors between image pairs.
CFGS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GSHT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fern Flower Orchids Room Trex Horns
Figure 4: Trajectory comparison on the LLFF dataset. Each row represents a method (CFGS, GSHT and Ours), and each column represents a scene.
TABLE IV: Pose estimation performance comparison on LLFF dataset.
@@ IV-B Evaluation Metrics
IV-C Implementation Details
We initialized camera poses and sparse Gaussian points using only scene images and camera intrinsics. During 3D Gaussian reconstruction, we alternately optimized 3D Gaussian points and camera poses. Global pose optimization was performed every 100 iterations, limited to the first 15,000 iterations. This restricted optimization strategy prevents error accumulation, as pose estimation errors could degrade reconstruction quality, which might further corrupt pose estimation accuracy. Thus, optimizing poses only during the initial quarter of training iterations is empirically justified. All experiments were conducted on a single RTX 3090 GPU. Unless otherwise stated, our experiments follow the same 3DGS parameter settings.
CFGS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GSHT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ballroom Barn Church Family Francis Horse
Figure 5: Trajectory comparison on the Tanks and Temples dataset. Each row represents a method (CFGS, GSHT and Ours), and each column represents a different scene.
@@ IV-C Implementation Details
TABLE V: Pose estimation performance comparison on Tanks and Temples dataset.
@@
IV-D Comparing with Baseline
Our experimental framework is built upon the original 3DGS architecture [1]. While the proposed modules are theoretically compatible with advanced 3DGS variants, our current implementation specifically adheres to the canonical formulation due to two methodological considerations: (1) make sure that the comparison with the original 3DGS can be made directly so that future generations can easily reproduce our work; and (2) isolating the performance impact of our contributions from other confounding factors. To maintain consistency, all architectural parameters strictly follow the original 3DGS configuration. This design choice facilitates direct comparability with COLMAP-based 3DGS baselines under identical experimental protocols.
TABLE VI: Pose estimation performance comparison between our method and VGGT on the LLFF dataset.
Scene RPEtrans RPErot ATE ↓
VGGT Ours VGGT Ours VGGT Ours
Fern 0.285 0.146 0.071 0.039 0.038 0.014
Flowers 0.160 0.100 0.036 0.052 0.004 0.005
Horns 0.056 0.051 0.009 0.027 0.010 0.019
Orchids 0.312 0.135 0.087 0.117 0.025 0.018
Room 0.104 0.039 0.035 0.030 0.017 0.005
Trex 0.138 0.084 0.032 0.026 0.011 0.006
Mean 0.203 0.093 0.059 0.049 0.017 0.011
IV-E Novel View Synthesis Evaluation
As shown in Tab. I, II and III, both CFGS and GSHT suffer degraded reconstruction quality, primarily due to their reliance on temporal continuity—when pose changes become large (e.g., under frame-subsampling on Tanks and Temples), their performance deteriorates sharply, as shown by the sharp drop in PSNR. In contrast, our method is sequence-agnostic and remains robust even under aggressive subsampling. As illustrated in Fig. 2, our method generates sharper geometric features and more coherent textures, in contrast to the blurred reconstructions of CFGS and the fragmented surfaces of GSHT. Beyond holistic visual assessment, Fig. 3 presents fine-grained comparisons of structural details. Our method achieves superior fidelity in geometric preservation and texture reconstruction compared to baseline approaches.
In addition, it is noteworthy that in Fig. 2, the CFGS method produces noticeably blurred novel view synthesis images due to its inaccurate pose estimation. This issue becomes particularly pronounced in the detailed regions as illustrated in Fig. 3. As demonstrated in the figure, both CFGS and GSHT exhibitscale drift and spatial misalignment of the display units. As shown in Fig. 3 (the first row), 3DGS shows obvious blurring around high-frequency structures such as the edge of the display. This is primarily due to the lack of joint camera pose optimization during training, where even slight pose inaccuracies can be amplified during dense rendering, leading to structural blur and color artifacts. We evaluate on the Shiny dataset, which features strong reflections and refractions. As shown in Tab. III, JOGS matches COLMAP+3DGS in overall metrics and significantly outperforms both COLMAP-free baselines.
IV-F Camera Pose Estimation
In Tab. IV and Tab. V, we provide a quantitative comparison of camera pose estimation performance on the LLFF and Tanks and Temples datasets, respectively. Following the alignment strategy proposed in GSHT, the estimated camera trajectories are first aligned with the ground truth in scale before evaluation. We report the results in terms of ATE and RPE. As shown in these results, our method achieves the lowest errors on most scenes. Qualitative results are visualized in Fig. 4 and Fig. 5. Our method produces estimated trajectories that are significantly more consistent with the ground truth than the baselines, demonstrating its robustness and effectiveness in handling complex camera motions and large-scale environments.
Furthermore, we compare our method with VGGT [50], a feed-forward neural network that reports strong performance across multiple 3D tasks, including camera parameter estimation. As shown in Table VI, on the LLFF dataset, our method achieves lower ATE and RPE on most scenes, with a lower mean error overall.
IV-G Ablation Study
To validate the necessity of our joint optimization framework, we conduct an ablation study comparing two variants: (1) Initialization-only (using initialized poses without iterative refinement during training) and (2) Full method (with alternating Gaussian and pose optimization).
As shown in Tab. VII, the full method outperforms the reduced initialization-only variant across all datasets. Init means working with only pose initialization without iterative refinement, while Full means working with the full version of our method containing joint optimization of Gaussian points and camera poses. The ablation study demonstrates that our combined optimization framework effectively mitigates error accumulation and enhances the synthesis accuracy of novel view scenes by alternately updating Gaussian points and refining camera poses using LK3D.
TABLE VII: Ablation study of joint optimization across three benchmark datasets.
[…]
Init Full Init Full Init Full
LLFF-NeRF 25.25 25.39 0.81 0.81 0.20 0.20
Tanks and Temples 25.94 26.91 0.86 0.88 0.14 0.13
Shiny 24.98 25.58 0.80 0.85 0.23 0.21
Mean 25.39 25.96 0.82 0.85 0.19 0.18
V Conclusion
In this paper, we introduce a novel view synthesis framework that jointly optimize pose estimation and 3DGS, without requiring camera poses as inputs. This framework outperforms state-of-the-art methods in both pose estimation accuracy and rendering quality, particularly under challenging conditions, by leveraging an alternating optimization strategy for 3D Gaussian representations and camera poses.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62376020).